# Asymptotically (un)safe scattering amplitudes from scratch: a deep dive into the IR jungle

### v1

@@ 2 Scalar scattering amplitudes
In this section, we set the stage and discuss the concrete amplitude that we want to compute in general terms. Specifically, we are interested in two shift-symmetric scalar fields $\phi$ and $\chi$ that are coupled to gravity. Before going into the details, let us briefly discuss the status of symmetries in quantum gravity, as it is relevant to motivate why we are interested in shift-symmetric scalars. There is a widely-believed conjecture that quantum gravity forbids the existence of global symmetries. More precisely, any would-be global symmetry ought to be gauged, or experience Planck-suppressed violations induced by quantum gravity. This conjecture is based on semi-classical black hole arguments and holds in string theory Banks:1988yz ; Giddings:1987cg ; Lee:1988ge ; Abbott:1989jw ; Coleman:1989zu ; Kamionkowski:1992mf ; Holman:1992us ; Kallosh:1995hi ; Banks:2010zn . It also motivates many of the swampland conjectures Vafa:2005ui ; Ooguri:2006in ; Brennan:2017rbf ; Palti:2019pca ; vanBeest:2021lhn ; Grana:2021zvf ; Agmon:2022thq . Nevertheless, there are known examples in which the conjecture does not hold Harlow:2020bee . In Asymptotic Safety all results to date Eichhorn:2011pc ; Eichhorn:2012va ; Labus:2015ska ; Percacci:2015wwa ; Eichhorn:2017eht ; Eichhorn:2020sbo ; Ali:2020znq ; deBrito:2021pyi ; Eichhorn:2021qet ; Laporte:2021kyp ; Eichhorn:2025ilu ; Assant:2025gto (see also Eichhorn:2022gku for an overview) point towards that global symmetries, including shift symmetry, are not broken by the RG flow if the regularisation does not break them artificially. Nevertheless, these results are based on approximations, in particular they are derived in Euclidean signature. It is conceivable that a Lorentzian treatment that properly includes black hole configurations could yield different results. Another option is that Asymptotic Safety gives rise to non-standard black hole thermodynamics Basile:2025zjc , thereby avoiding the semi-classical arguments. Lastly, there could also be a dynamical suppression of black hole configurations in the path integral Borissova:2020knn ; Borissova:2024hkc .
@@
Approximating the wave function renormalisations of the scalars, $Z_{\phi}$ and $Z_{\chi}$, by unity has been shown to be an excellent approximation in momentum-dependent computations Meibohm:2015twa ; Knorr:2019atm ; Kher:2025rve , see also e.g. Dona:2013qba ; Dona:2015tnf ; Eichhorn:2016esv ; Eichhorn:2017eht ; Eichhorn:2017sok ; Eichhorn:2017als ; Eichhorn:2020sbo ; deBrito:2021pyi ; Laporte:2021kyp ; Pastor-Gutierrez:2022nki ; deBrito:2025nog for works computing a constant anomalous dimension, and Henz:2016aoh for a field-dependent wave function. The gravitational form factors $F_{R}$, $F_{Ric}$ have been computed in various approximations Christiansen:2012rx ; Codello:2013fpa ; Christiansen:2014raa ; Bosma:2019aiu ; Bonanno:2021squ ; Knorr:2021niv ; Fehre:2021eob ; Knorr:2023usb ; Pastor-Gutierrez:2024sbt ; Kher:2025rve ; Pawlowski:2025etp and display a non-trivial behaviour which we however neglect to be able to treat the system analytically. [^1] In our simple approximation, the inclusion of these form factors can largely be compensated by a different regulator choice, see the final flow equation (23) below. The only clear exception is the large-momentum behaviour of the propagator, which will be modified under the inclusion of these form factors, in turn adding an anomalous dimension to the form factor that we resolve. The same difference in scaling will also affect the mediated diagram of the scattering amplitude (left panel of Figure 1), so that we expect the overall picture to not change qualitatively. By contrast, the non-minimal interactions between gravity and shift-symmetric scalar fields have only been computed in local approximations so far Eichhorn:2017sok ; Laporte:2021kyp ; Knorr:2022ilz ; Ohta:2025xxo ; deBrito:2025nog , so no information on the quality of such an approximation is available yet. See also e.g. Narain:2009fy ; Henz:2013oxa ; Percacci:2015wwa ; Labus:2015ska ; Henz:2016aoh ; Pawlowski:2018ixd ; Wetterich:2019zdo ; Eichhorn:2020sbo ; Sen:2021ffc ; Maitiniyazi:2025pou for non-minimal interactions between gravity and scalars without shift symmetry.
@@ 4.1 Momentum-dependent results
$I(z)=\frac{2}{\pi^{2}}\left[\frac{1-e^{-\frac{z}{2}}}{z}-\frac{\eta_{N,t}}{4}\left(\frac{1+3e^{-\frac{2z}{3}}-4e^{-\frac{z}{2}}}{z}+2\left(\operatorname{Ei}(-\tfrac{2z}{3})-\operatorname{Ei}(-\tfrac{z}{2})\right)\right)\right]\,.$ (28)
In this expression, $\operatorname{Ei}$ denotes the exponential integral function.
@@
$F^{E}(P^{2})=128G_{N}^{2}\,e^{\frac{G_{N}P^{2}}{2g_{\ast}}}\,\operatorname{Ei}(-\tfrac{G_{N}P^{2}}{2g_{\ast}})\,.$ (35)
@@
Since we have the analytic expression for the form factor, we can perform the Wick rotation and evaluate it for squared momenta $P^{2}\in\mathbbm{R}$. [^11] There is a branch cut on the positive real line, which can give rise to an imaginary part based on the choice of Riemann sheet. We will not discuss this here, and focus on the real part only. We show the form factor in Figure 4.
Figure 4: Euclidean form factor $F^{E}$ in Planck units, as a function of Euclidean squared momenta in terms of the quantum gravity scale $G_{N}/2g_{\ast}$. Since we have a completely analytical expression, we can evaluate the form factor also for negative arguments. At vanishing argument, there is a logarithmic divergence, see (40).
@@
For large (negative or positive) momentum, the form factor falls off,
$F^{E}(P^{2})\sim-\frac{256g_{\ast}G_{N}}{P^{2}}\,,\qquad P^{2}\to\pm\infty\,.$ (41)
@@
$P\partial_{P}F^{E}(P^{2})=256G_{N}^{2}\left(1+\frac{G_{N}P^{2}}{2g_{\ast}}\,e^{\frac{G_{N}P^{2}}{2g_{\ast}}}\operatorname{Ei}(-\tfrac{G_{N}P^{2}}{2g_{\ast}})\right)=256G_{N}^{2}+2\frac{G_{N}P^{2}}{2g_{\ast}}\,F^{E}(P^{2})\,.$ (42)
We depict it in Figure 5. We can see that for large (negative or positive) $P^{2}$, this beta function goes to zero. In this sense, the form factor displays a fixed point at large momentum.
Figure 5: Momentum beta function in Planck units, as a function of Euclidean momenta in terms of the quantum gravity scale. Since we have a completely analytical expression, we can evaluate the form factor also for negative arguments. At vanishing argument, the function is continuous but not differentiable due to the logarithm in the form factor, cf. (40) and (42).
@@
$F_{1}(s)=-F_{2}(s)=128G_{N}^{2}e^{-\frac{G_{N}s}{2g_{\ast}}}\operatorname{Ei}(\tfrac{G_{N}s}{2g_{\ast}})\,.$ (44)
Using our general formula for the partial wave amplitudes, we find at large energies $s\to\infty$,
@@
We show the partial wave amplitudes in comparison to their tree-level counterparts in Figure 6, for two choices of $g_{\ast}$. In the left panel, we choose $g_{\ast}=\pi/4$ which bounds $a_{0}$ for all energies. In the right panel, we choose $g_{\ast}=\pi/2$, for which both $a_{0}$ and $a_{2}$ are unbounded. Both choices of $g_{\ast}$ are broadly compatible with the literature. We find that for these choices, the partial wave amplitudes violate the unitarity bound at lower energies than at tree level – except of course for $a_{0}$ and $g_{\ast}=\pi/4$.
Figure 6: Comparison of partial wave amplitudes with their tree-level counterparts as a function of energy transfer in Planck units, for two choices of the fixed point value $g_{\ast}$. In the left panel, $g_{\ast}=\pi/4$ so that $a_{0}$ is bounded, but $a_{2}$ is not bounded and exceeds unit strength before the tree-level amplitude does. In the right panel, $g_{\ast}=\pi/2$, and both $a_{0}$ and $a_{2}$ exceed unity before their tree-level counterparts.
@@ 7 Conclusion and outlook
To gain a more reliable understanding of scattering processes in Asymptotic Safety, it is clearly mandatory to improve the approximation and to discuss other matter content that is more realistic. While the momentum dependence of propagators is by now under control Christiansen:2012rx ; Codello:2013fpa ; Christiansen:2014raa ; Meibohm:2015twa ; Knorr:2019atm ; Bosma:2019aiu ; Bonanno:2021squ ; Knorr:2021niv ; Fehre:2021eob ; Knorr:2023usb ; Pastor-Gutierrez:2024sbt ; Kher:2025rve ; Pawlowski:2025etp , less is known about interaction vertices Christiansen:2015rva ; Denz:2016qks ; Christiansen:2017cxa ; Knorr:2017fus ; Eichhorn:2018akn ; Eichhorn:2018ydy ; Eichhorn:2018nda ; deBrito:2025nog . Resolving their momentum dependence is crucial not only to establish physical Asymptotic Safety, but also to understand whether the theory is unitary and causal. Moreover, matter fluctuations have to be taken into account. Another restriction that needs to be lifted is that of using a flat background. It seems reasonable to expect that some of the problems of massless theories disappear when curvature dependence is included, whenever the curvature acts as an effective mass for the massless fluctuations.
@@ Appendix A NLO
$\qquad-\frac{128g_{\ast}^{2}(G_{N}k^{2})^{2}}{(g_{\ast}+G_{N}k^{2})^{2}}\left(\operatorname{Ei}(-\tfrac{2z}{3})-\operatorname{Ei}(-\tfrac{z}{2})\right)$
@@
$F^{E,\text{NLO}}(P^{2})=128G_{N}^{2}e^{\frac{G_{N}P^{2}}{2g_{\ast}}}\left(2\operatorname{Ei}(-\tfrac{G_{N}P^{2}}{2g_{\ast}})-e^{\frac{G_{N}P^{2}}{6g_{\ast}}}\operatorname{Ei}(-\tfrac{2G_{N}P^{2}}{3g_{\ast}})\right)\,.$ (83)
We compare the leading-order with the next-to-leading order result in Figure 9. We find that qualitatively, both form factors agree. Both have the same scaling for small and large momenta, and only differ in the numerical prefactors. Interestingly, we can write
@@
where $F^{E,\text{LO}}$ is the leading-order form factor as given in the main text, (35). This relation makes it straightforward to perform the same analysis about small and large momentum limits, so we will not spell out the details. When once again comparing to the EFT, we can extract the refined cutoff scale
@@
Figure 9: Comparison of the Euclidean form factor at leading and next-to-leading order. Since we have a completely analytical expression, we can evaluate the form factor also for negative arguments. The qualitative momentum dependence is the same, and only small quantitative differences appear.

### v2

@@ 2 Scalar scattering amplitudes
In this section, we set the stage and discuss the concrete amplitude that we want to compute in general terms. Specifically, we are interested in two shift-symmetric scalar fields $\phi$ and $\chi$ that are coupled to gravity. Before going into the details, let us briefly discuss the status of symmetries in quantum gravity, as it is relevant to motivate why we are interested in shift-symmetric scalars. There is a widely-believed conjecture that quantum gravity forbids the existence of global symmetries. More precisely, any would-be global symmetry ought to be gauged, or experience Planck-suppressed violations induced by quantum gravity. This conjecture is based on semi-classical black hole arguments and holds in string theory Banks:1988yz ; Giddings:1987cg ; Lee:1988ge ; Abbott:1989jw ; Coleman:1989zu ; Kamionkowski:1992mf ; Holman:1992us ; Kallosh:1995hi ; Banks:2010zn . It also motivates many of the swampland conjectures Vafa:2005ui ; Ooguri:2006in ; Brennan:2017rbf ; Palti:2019pca ; vanBeest:2021lhn ; Grana:2021zvf ; Agmon:2022thq . Nevertheless, there are examples in which the conjecture does not hold Harlow:2020bee (see, however, Geng:2025gns ). In Asymptotic Safety all results to date Eichhorn:2011pc ; Eichhorn:2012va ; Labus:2015ska ; Percacci:2015wwa ; Eichhorn:2017eht ; Eichhorn:2020sbo ; Ali:2020znq ; deBrito:2021pyi ; Eichhorn:2021qet ; Laporte:2021kyp ; Eichhorn:2025ilu ; Assant:2025gto (see also Eichhorn:2022gku for an overview) point towards that global symmetries, including shift symmetry, are not broken by the RG flow if the regularisation does not break them artificially. Nevertheless, these results are based on approximations, in particular they are derived in Euclidean signature. It is conceivable that a Lorentzian treatment that properly includes black hole configurations could yield different results. Another option is that Asymptotic Safety gives rise to non-standard black hole thermodynamics Basile:2025zjc , thereby avoiding the semi-classical arguments. Lastly, there could also be a dynamical suppression of black hole configurations in the path integral Borissova:2020knn ; Borissova:2024hkc .
@@
Approximating the wave function renormalisations of the scalars, $Z_{\phi}$ and $Z_{\chi}$, by unity has been shown to be an excellent approximation in momentum-dependent computations Meibohm:2015twa ; Knorr:2019atm ; Kher:2025rve , see also e.g. Dona:2013qba ; Dona:2015tnf ; Eichhorn:2016esv ; Eichhorn:2017eht ; Eichhorn:2017sok ; Eichhorn:2017als ; Eichhorn:2020sbo ; deBrito:2021pyi ; Laporte:2021kyp ; Pastor-Gutierrez:2022nki ; deBrito:2025nog for works computing a constant anomalous dimension, and Henz:2016aoh for a field-dependent wave function. The gravitational form factors $F_{R}$, $F_{Ric}$ have been computed in various approximations Christiansen:2012rx ; Codello:2013fpa ; Christiansen:2014raa ; Bosma:2019aiu ; Bonanno:2021squ ; Knorr:2021niv ; Fehre:2021eob ; Knorr:2023usb ; Pastor-Gutierrez:2024sbt ; Kher:2025rve ; Pawlowski:2025etp and display a non-trivial behaviour which we however neglect to be able to treat the system analytically. [^1] In our simple approximation, the inclusion of these form factors can largely be compensated by a different regulator choice, see the final flow equation (23) below. The only clear exception is the large-momentum behaviour of the propagator, which will be modified under the inclusion of these form factors, in turn adding an anomalous dimension to the form factor that we resolve. The same difference in scaling will also affect the mediated diagram of the scattering amplitude (left panel of Figure 1), so that we expect the overall picture to not change qualitatively. By contrast, the non-minimal interactions between gravity and shift-symmetric scalar fields have mostly been computed in local approximations Eichhorn:2017sok ; Laporte:2021kyp ; Knorr:2022ilz ; Ohta:2025xxo ; deBrito:2025nog , see however Chiesa:2026tlz for recent work on its momentum dependence. See also e.g. Narain:2009fy ; Henz:2013oxa ; Percacci:2015wwa ; Labus:2015ska ; Henz:2016aoh ; Pawlowski:2018ixd ; Wetterich:2019zdo ; Eichhorn:2020sbo ; Sen:2021ffc ; Maitiniyazi:2025pou for non-minimal interactions between gravity and scalars without shift symmetry.
@@ 4.1 Momentum-dependent results
$I(z)=\frac{2}{\pi^{2}}\left[\frac{1-e^{-\frac{z}{2}}}{z}-\frac{\eta_{N,t}}{4}\left(\frac{1+3e^{-\frac{2z}{3}}-4e^{-\frac{z}{2}}}{z}-2\left(\Gamma(0,\tfrac{2z}{3})-\Gamma(0,\tfrac{z}{2})\right)\right)\right]\,.$ (28)
In this expression, $\Gamma(a,b)$ is the incomplete Gamma function.
@@
$F^{E}(P^{2})=-128G_{N}^{2}\,e^{\frac{G_{N}P^{2}}{2g_{\ast}}}\,\Gamma(0,\tfrac{G_{N}P^{2}}{2g_{\ast}})\,.$ (35)
@@
The form factor is analytic everywhere in the complex plane, except for a branch cut on the negative real axis. This is expected as it corresponds to the multi-particle continuum. In particular, along the cut, we have
$\text{Im}\,F^{E}(P^{2})=128\pi G_{N}^{2}\,e^{\frac{G_{N}P^{2}}{2g_{\ast}}}\,,\qquad P^{2}\leq 0\,.$ (40)
This allows us to perform the Wick rotation and evaluate the form factor for squared momenta $P^{2}\in\mathbbm{C}$. We show the form factor on the real $P^{2}$-axis in Figure 4, and in the complex plane in Figure 5.
Figure 4: Real and imaginary parts of the Euclidean form factor $F^{E}$ in Planck units, as a function of Euclidean squared momenta in terms of the quantum gravity scale $G_{N}/2g_{\ast}$. At vanishing argument, there is a logarithmic divergence in the real part, see (41).
@@
Figure 5: Euclidean form factor $F^{E}$ in Planck units, as a function of complex Euclidean squared momenta in terms of the quantum gravity scale $G_{N}/2g_{\ast}$. The colour gradient indicates the phase of the function, and the dashed line indicates the branch cut.
For large (complex) momentum, the form factor falls off,
$F^{E}(P^{2})\sim-\frac{256g_{\ast}G_{N}}{P^{2}}\,,\qquad|P^{2}|\to\infty\,,P^{2}\notin\mathbbm{R}_{-}\,.$ (42)
@@
$P\partial_{P}F^{E}(P^{2})=256G_{N}^{2}\left(1-\frac{G_{N}P^{2}}{2g_{\ast}}\,e^{\frac{G_{N}P^{2}}{2g_{\ast}}}\Gamma(0,\tfrac{G_{N}P^{2}}{2g_{\ast}})\right)=256G_{N}^{2}+2\frac{G_{N}P^{2}}{2g_{\ast}}\,F^{E}(P^{2})\,.$ (43)
We depict its real part in Figure 6. We can see that for large (negative or positive) $P^{2}$, this beta function goes to zero. In this sense, the form factor displays a fixed point at large momentum.
Figure 6: Real part of the momentum beta function in Planck units, as a function of Euclidean momenta in terms of the quantum gravity scale. Since we have a completely analytical expression, we can evaluate the form factor also for negative arguments. At vanishing argument, the function is continuous but not differentiable due to the logarithm in the form factor, cf. (41) and (43).
@@
$F_{1}(s)=-F_{2}(s)=-128G_{N}^{2}e^{-\frac{G_{N}s}{2g_{\ast}}}\Gamma(0,-\tfrac{G_{N}s}{2g_{\ast}})\,.$ (45)
Using our general formula for the partial wave amplitudes, we find at large energies $s\to\infty$, [^11] Since the imaginary part falls off exponentially, it plays no role for our considerations here.
@@
We show the real parts of the partial wave amplitudes in comparison to their tree-level counterparts in Figure 7, for two choices of $g_{\ast}$. In the left panel, we choose $g_{\ast}=\pi/4$ which bounds $a_{0}$ for all energies. In the right panel, we choose $g_{\ast}=\pi/2$, for which both $a_{0}$ and $a_{2}$ are unbounded. Both choices of $g_{\ast}$ are broadly compatible with the literature. We find that for these choices, the partial wave amplitudes violate the unitarity bound at lower energies than at tree level – except of course for $a_{0}$ and $g_{\ast}=\pi/4$.
Figure 7: Comparison of real parts of the partial wave amplitudes with their tree-level counterparts as a function of energy transfer in Planck units, for two choices of the fixed point value $g_{\ast}$. In the left panel, $g_{\ast}=\pi/4$ so that $a_{0}$ is bounded, but $a_{2}$ is not bounded and exceeds unit strength before the tree-level amplitude does. In the right panel, $g_{\ast}=\pi/2$, and both $a_{0}$ and $a_{2}$ exceed unity before their tree-level counterparts.
@@ 5.2 Momentum-dependent results with masses
As in the massless case, (71) is analytic in the entire complex plane except for the negative real line, where it admits a branch cut. Along this cut,
[…]
$-6P^{2}\left(m_{\phi,\ast}^{2}M_{\chi}^{2}+m_{\chi,\ast}^{2}M_{\phi}^{2}\right)\,{}_{1}F_{1}\left(2+\frac{5g_{\ast}}{\pi},3,\frac{G_{N}P^{2}}{2g_{\ast}}\right)$
$+P^{4}m_{\phi,\ast}^{2}m_{\chi,\ast}^{2}\,{}_{1}F_{1}\left(2+\frac{5g_{\ast}}{\pi},4,\frac{G_{N}P^{2}}{2g_{\ast}}\right)\Bigg]\,,\qquad P^{2}\leq 0\,.$
@@ 7 Conclusion and outlook
To gain a more reliable understanding of scattering processes in Asymptotic Safety, it is clearly mandatory to improve the approximation and to discuss other matter content that is more realistic. While the momentum dependence of propagators is by now under control Christiansen:2012rx ; Codello:2013fpa ; Christiansen:2014raa ; Meibohm:2015twa ; Knorr:2019atm ; Bosma:2019aiu ; Bonanno:2021squ ; Knorr:2021niv ; Fehre:2021eob ; Knorr:2023usb ; Pastor-Gutierrez:2024sbt ; Kher:2025rve ; Pawlowski:2025etp , less is known about interaction vertices Christiansen:2015rva ; Denz:2016qks ; Christiansen:2017cxa ; Knorr:2017fus ; Eichhorn:2018akn ; Eichhorn:2018ydy ; Eichhorn:2018nda ; deBrito:2025nog , for recent progress on the scalar-scalar-graviton vertex see Chiesa:2026tlz . Resolving their momentum dependence is crucial not only to establish physical Asymptotic Safety, but also to understand whether the theory is unitary and causal. Moreover, matter fluctuations have to be taken into account. Another restriction that needs to be lifted is that of using a flat background. It seems reasonable to expect that some of the problems of massless theories disappear when curvature dependence is included, whenever the curvature acts as an effective mass for the massless fluctuations.
@@ Appendix A NLO
$\qquad+\frac{128g_{\ast}^{2}(G_{N}k^{2})^{2}}{(g_{\ast}+G_{N}k^{2})^{2}}\left(\Gamma(0,\tfrac{2z}{3})-\Gamma(0,\tfrac{z}{2})\right)$
@@
$F^{E,\text{NLO}}(P^{2})=-128G_{N}^{2}e^{\frac{G_{N}P^{2}}{2g_{\ast}}}\left(2\Gamma(0,\tfrac{G_{N}P^{2}}{2g_{\ast}})-e^{\frac{G_{N}P^{2}}{6g_{\ast}}}\Gamma(0,\tfrac{2G_{N}P^{2}}{3g_{\ast}})\right)\,.$ (85)
We compare the real parts of the leading-order with the next-to-leading order result in Figure 10. We find that qualitatively, both form factors agree. Both have the same scaling for small and large momenta, and only differ in the numerical prefactors. Interestingly, we can write
@@
where $F^{E,\text{LO}}$ is the leading-order form factor as given in the main text, (35). This relation makes it straightforward to perform the same analysis about small and large momentum limits as well as the branch cut, so we will not spell out the details. When once again comparing to the EFT, we can extract the refined cutoff scale
@@
Figure 10: Comparison of the real parts of the Euclidean form factor at leading and next-to-leading order. Since we have a completely analytical expression, we can evaluate the form factor also for negative arguments. The qualitative momentum dependence is the same, and only small quantitative differences appear.